OpenStudy (anonymous):

Find all solutions to the given equation in the interval [0,2pi] (sin x + 1)(tan x +1 ) = 0

OpenStudy (anonymous):

i m writing in general form if (x + a) (x + b) = 0 x+a = 0 or x+b = 0 hope this will help

OpenStudy (anonymous):

So sin(x)=-1 and tan(x)=-1

OpenStudy (anonymous):

now give the solution in radians or degrees keeping in mind the [0,2pi]

OpenStudy (anonymous):

That's the part I'm most confused on. It has to be in radians but how do I find it? I know it's on the unit circle but I don't remember.

OpenStudy (anonymous):

 in first quadrant ( 0< theta ≤ 90 ) all sine cosine and tangent are positive in second quadrant ( 90< theta ≤ 180) only sine is positive in third quadrant (180< theta ≤270) only tangent is positive in fouth quadrant (270< theta ≤360) only cosine is positive